Explain how opposites are applied when solving one-step equations. How do you know what to use to solve a one-step equation? Be specific and complete.

Answer by Theo(13342)

here's a good reference.

http://www.algebralab.org/lessons/lesson.aspx?file=algebra_onevariableonestep.xml

a one step equation is solved in one step, as its name implies.

the goal is to get the variable on one sides of the equation and the constant on the other side.

the basic rules are:

whatever you do to one side of the equation you have to do to the aother side in order to preserve the equality.

you use opposites to move constants and or coefficients from one side of the equation to the other.

some examples:

-----------------------------------------

x + 5 = 3

subtract 5 (the opposite of adding 5) from both sides of the equation to get:

x + 5 - 5 = 3 - 5

combine like terms to get x = -2

-------------------------------------------

x - 3 = 5

add 3 (the opposite of subtracting 3) to both sides of the equation to get:

x - 3 + 3 = 5 + 3

combine like terms to get x = 8

-----------------------------------------

5 * x = 15

divide both sides of the equation by 5 (the opposite of multiplying by 5) to get:

5 * x / 5 = 15 / 5

simplify to get x = 3

-------------------------------------

x / 20 = 3

multiply both sides of the equation by 20 (the opposited of dividing by 20) to get:

x / 20 * 20 = 3 * 20

simplify to get x = 60

the reference has more examples.

check them out.